# Learning Network Representations with Disentangled Graph Auto-Encoder

### v1

@@ 3 Preliminaries and Problem Formulation
In order to better describe the issues that our article focuses on and introduce our work, we first briefly review the idea of GAE and VGAE in section 3.1 and then formulate the problems we are concerned about in section 3.2.
3.1 Preliminaries on (Vriational) Graph Auto-Encoder
@@
VGAE and GAE introduced by Kipf et al. (2016a) are frameworks designed for unsupervised learning on graph-structured data. One notable feature is their natural ability to incorporate node features, leading to a substantial improvement in predictive performance on different tasks across various benchmark datasets.
@@
GCN is used as the encoder in Kipf et al. (2016a), which is designed specifically for processing and extracting information from graph-structured data. For a $L$-layer GCN, its message propagation rule is defined by the following formula:
@@
Here, $\theta$ is the parameters of the encoder, $\tilde{A}=A+I_{N}$, $I_{N}$ is the identity matrix, $\tilde{D}_{ii}=\sum_{j=1}^{N}\tilde{A}_{ij}$ and $W^{(l)}$ is the parameters of weight matrix. ${\rm Relu}(\cdot)$ denotes the Relu activation function. $H^{(l)}$ is the matrix of activations in the $l$-th layer, $H^{(0)}=X$ and $H^{(L)}=Z$. In GAE, Kipf et al. (2016a) used two-layer GCN as the encoder:
@@
where $\mu_{i}$ and ${\rm log}\,\sigma_{i}^{2}$ are learned from two GCNs with shared first-layer training parameters. Samples of $q(Z|X,A)$ can be obtained by using the reparameterization trick (Doersch, 2016).
The decoder is based on the inner product between latent representations of each pair of nodes. In the case of GAE, the reconstruction of the adjacency matrix ${A}$ is formulated as follows:
@@
where $\sigma$ is typically the sigmoid function. For VGAE, the decoder is probabilistic and is defined by the inner product between latent variables. This can be expressed as follows:
@@
The training objective for VGAE is to recover information about the graph from the graph embedding matrix $Z$. Therefore, our aim is to learn the model parameters by maximizing the log marginal distribution which can be lower-bounded by ELBO:
@@
$=$ $\mathbb{E}_{Z\sim q_{\theta}(Z|X,A)}\log p(A|Z)+\mathbb{E}_{Z\sim q_{\theta}(Z|X,A)}\log\frac{p(Z)}{q_{\theta}(Z|X,A)}$ (10)
$=$ $\mathbb{E}_{Z\sim q_{\theta}(Z|X,A)}\log p(A|Z)-D_{KL}(q_{\theta}(Z|X,A)||p(Z)).$ (11)
Here, the inequality is derived from Jensen’s inequality (Weisstein, 2006), and $D_{KL}(\cdot||\cdot)$ represents the Kullback-Leibler (KL) divergence (Kullback et al., 1951). We set $p(Z)$ to be a Gaussian normal distribution, i.e.,
@@
3.2 Problem Formulation
For (variational) graph auto-encoder, the objective is to learn a graph encoder $f_{\theta}(\cdot)$ with parameters $\theta$ that outputs interpretable low-dimensional stochastic latent variables $Z=f_{\theta}(G)\in\mathbb{R}^{N\times d}$ ($d\ll n$). Simultaneously, such learned node representations, similar to other node representation learning methods, are capable of tackling a diverse range of tasks related to nodes. Here, our primary focus is on the classical task of link prediction for nodes, aligning with the experiments conducted in Kipf et al. (2016a).
In our study, we aim to improve the link prediction performance in (variational) graph auto-encoder by improving its overall framework, with the primary goal of learning disentangled node representations through the disentangled graph encoder $f_{\theta}$. For each node $u$, the output $z_{u}$ from the disentangled encoder will be a disentangled representation. Specifically, assuming there are $K$ latent factors to be disentangled, $z_{u}$ is composed of $K$ independent components, i.e., $z_{u}=\left[z_{u,1},z_{u,2},\dots,z_{u,K}\right]$, where $z_{i,k}\in\mathbb{R}^{\frac{d}{K}}$ ($1\leq k\leq K$). The $k$-th component $z_{u,k}$ is used to characterize aspects of node $u$ that are related to the $k$-the latent factor.
Here, we assume the underlying factors generating observed graph-structured data are mutually independent, which is a common premise in the field of disentangled representation learning (Higgins et al., 2018; Bengio et al., 2013). Our goal is to learn factorized and expressive representations where each component is independent and exclusively corresponds to a single ground-truth factor behind the graph formulation. This is intended to enhance the completeness of each node’s representation, resulting in more effective node descriptions and improved performance in graph analytic tasks.
4 Proposed Mechanism
@@
4.1 Dynamic Disentangled Graph Encoder
Here, our goal is to introduce the disentangled graph encoder that can learn disentangled node embeddings. Essentially, it comprises the disentangled graph convolutional neural network, which we refer to as DGCN. We assume that each node consists of $K$ independent components, indicating the presence of K latent factors that need to be disentangled. For a node $u\in V$, its hidden representation $z_{u}$ is expressed as $\left[z_{u,1},z_{u,2},\dots,z_{u,K}\right]$, where $z_{u,k}\in\mathbb{R}^{\frac{d}{K}}$ ($1\leq k\leq K$) is used to describe $k$-th aspects of node $u$.
@@ 4.1 Dynamic Disentangled Graph Encoder
where $W_{k}\in\mathbb{R}^{n\times\frac{d}{K}}$ and $b_{k}\in\mathbb{R}^{\frac{d}{K}}$ are the parameters of $k$-th subspace. For a comprehensive understanding of aspect $k$ for node $u$, we need to extract information from the neighborhood of $u$. This involves constructing $z_{u,k}$ using both $c_{u,k}$ and $\{c_{v,k}:(u,v)\in E\}$. Next, we provide three hypotheses for the relationships between nodes:
Hypothesis 1: The generation of links between nodes result from the interaction of various hidden factors.
Hypothesis 2: If there is similarity in the $k$-th component space between node $u$ and its neighboring node $v$, it indicates that the link between the two nodes may be attributed to factor $k$.
@@
Hypothesis 1 aligns with our realistic understanding of the reasons behind the formation of graphs. It suggests that the links between nodes can be influenced by various factors. In comparison to previous research on disentanglement, this hypothesis is closer to real-world situations (Liu et al., 2020; Ma et al., 2019). For example, in a friendship network, the acquaintance between two individuals may result from relationships such as colleagues, classmates, or shared interests. Hypotheses 2 and 3 correspond to the first-order and second-order proximities within the topological structure influenced by factor $k$, respectively (Tang et al., 2015). This insight also inspires us to explore $K$ subspaces to identify latent factors.
Building upon these three hypotheses, we learn the disentangled representations of nodes by leveraging dynamic assignment mechanism (Zheng et al., 2021), which is a disentangle layer:
@@
where iteration $t=1,2,\dots,T$. $p_{u,v}^{k}$ represents the probability of the connection between nodes $v$ and $u$ attributed to the $k$-th factor, while $q_{v,u}^{k}$ denotes the importance of aggregating node $v$ with respect to node $u$ in the $k$-th component space. They satisfies $p_{u,v}^{k}\geq 0$, $q_{v,u}^{k}\geq 0$, $\sum_{k=1}^{K}p_{u,v}^{k}=1$ and $\sum_{k=1}^{K}q_{v,u}^{k}=1$. $\alpha$ and $\beta$ serve as trade-off coefficients and we treat them as trainable parameters and adjust their range to be within $(0,1)$. The dynamic assignment mechanism will iteratively infer $p_{u,v}^{k}$ ,$q_{v,u}^{k}$ and construct $z_{u,k}$. Meanwhile, this mechanism enables the coexistence of multiple relationships between nodes. It’s important to note that there are a total of $L$ disentangle layers, and in each layer $l\leq L-1$, the value of $c_{u,k}$ is ultimately assigned as $z_{u,k}^{T}$. For additional details, please refer to Algorithm 1.
@@
Subsequently, by applying Eqs. (2), (3), (4) and replacing GCN with DGCN, we obtain the dynamic disentangled graph encoder.
In contrast to existing holistic graph encoders, the dynamic disentangled graph encoder comprises $K$ channels, enabling the identification of complex heterogeneous latent factors and the depiction of multiple relation aspects.
4.2 Enhancing Posterior Flexibility through Component-wise Normalizing Flows
We observe that when the inference process relies solely on the dynamic disentangled graph encoder, DVGA will typically output a factorized distribution. This implies that the individual component representation $z_{u,k}$ are not only independent of each other but also mutually independent for each $d$-dimensional vectors. The latter is neither necessary nor desired because we want each $d$-dimensional component to better capture enough information about each factor. Therefore, in this section, we aim to enhance the flexibility of each $d$-dimensional component representation.
Normalizing Flows (NF) (Papamakarios et al., 2017) are frequently used to enrich the families of posterior distribution. They impose explicit density functions for the mixing distributions in the hierarchy. Thus, it is sometimes employed in VGAE (Hasanzadeh et al., 2019; Shan et al., 2020) for the purpose of enhancing model flexibility. To better learn the information of each factor, we introduce a flow-based, component-wise distribution as the posterior distribution for the latent variables in DVGA. This distribution is designed to make the $d$-dimensional component latent variables, namely, embeddings derived from each channel, better at describing the complex factor properties.
Our main flow process involves applying separate flows to the $K$ components of $z_{u}$, with each flow consisting of $M$ steps (e.g., $M$ = 4). Each step incorporates an affine coupling layer, which primarily ensures that the latent variable $z_{u,k}$ follows a flexible distribution. Specifically, for each component $1\leq k\leq K$, we iterate as follows to apply component-wise flows, i.e., for $m=1,2,\dots,M$:
@@ 4.2 Enhancing Posterior Flexibility through Component-wise Normalizing Flows
Here, $s_{k}(\cdot)$ and $t_{k}(\cdot)$ denote transformation functions (mapping the input from $d^{\prime}$ to $d-d^{\prime}$ dimension space), $s(z_{u,k,1:d^{\prime}}^{(m-1)})=\left[s_{1}(z_{u,k,1:d^{\prime}}^{(m-1)}),s_{2}(z_{u,k,1:d^{\prime}}^{(m-1)}),\dots,s_{d-d^{\prime}}(z_{u,k,1:d^{\prime}}^{(m-1)})\right]^{\top}$, $t(z_{u,k,1:d^{\prime}}^{(m-1)})=\left[t_{1}(z_{u,k,1:d^{\prime}}^{(m-1)}),t_{2}(z_{u,k,1:d^{\prime}}^{(m-1)}),\dots,t_{d-d^{\prime}}(z_{u,k,1:d^{\prime}}^{(m-1)})\right]^{\top}$. The $\odot$ represents the Hada-mard product. The output of the affine coupling layer is obtained by concatenating $z_{u,k,1:d^{\prime}}^{(m)}$ and $z_{u,k,d^{\prime}+1:d}^{(m)}$. We swap or randomly shuffle the dimensions to enhance the mixing of information between two affine coupling layers.
@@
4.3 Factor-wise Decoder
Unlike the conventional decoder design in the (Variational) Graph Auto-Encoder and other approaches that broadly utilize graph information to design decoders (Hasanzadeh et al., 2019; Grover et al., 2019; Shi et al., 2020), both DGA and DVGA introduce a novel factor-wise decoder for predicting edge connections between nodes. This design encourages the decoder to integrate information at the factor level, promoting link prediction performance.
As we already know, the formation of real-world graphs is typically influenced by multiple latent heterogeneous factors. Thus, if any factor leads to the existence of a relationship between two nodes, an edge will connect them in the final result. Based on this observation, we modify the traditional decoder. The structure of the proposed decoder is also shown in Figure 2. Firstly, we perform cosine similarity calculation for each $K$ components between each pair of nodes $z_{u}=\left[z_{u,1},...,z_{u,K}\right]^{\top}$ and $z_{v}=\left[z_{v,1},...,z_{v,K}\right]^{\top}$. The output vector $A_{factor}$ is in $\mathbb{R}^{K\times N\times N}$ , corresponding to predicting the connection status of nodes in $K$ component spaces associated with $K$ factors. The result of each component space connection corresponds to an $N\times N$ matrix. This vector $A_{factor}$ is then fused by performing a max-pooling operation on the prediction of each two-node connection of the $K$ channels to obtain the final edge connection state. The entire block, denoted as $\mathcal{D}$, is ultimately merged with inner products constructed from $Z$. This extra inner product connection functions in a manner similar to the residual connections found in residual networks (He et al., 2016). Finally, the output of the factor-wise decoder is as follows:

### v2

@@ III Preliminaries and Problem Formulation
To better describe the issues that our paper focuses on and present our work, we first briefly review the concepts of GAE and VGAE in Section III-A and then formulate the problems in Section III-B.
III-A Preliminaries on (Vriational) Graph Auto-Encoder
@@
VGAE and GAE, introduced by [12], are specifically designed for unsupervised learning on graph-structured data. A remarkable characteristic is their natural ability to incorporate node features, resulting in a significant enhancement in predictive performance on diverse tasks across many benchmark datasets.
@@
GCN is used as the encoder in VGAE and GAE, which is designed specifically for processing and extracting information from graph-structured data. For a $L$-layer GCN, its message propagation rule is defined by the following formula:
@@
Here, $\theta$ is the parameters of the encoder, $\tilde{A}=A+I_{N}$, $I_{N}$ is the identity matrix, $\tilde{D}_{ii}=\sum_{j=1}^{N}\tilde{A}_{ij}$ and $W^{(l)}$ is the parameters of weight matrix. ${\rm Relu}(\cdot)$ denotes the Relu activation function. $H^{(l)}$ is the matrix of activations in the $l$-th layer, $H^{(0)}=X$ and $H^{(L)}=Z$. In GAE, [12] used two-layer GCN as the encoder:
@@
where $\mu_{i}$ and ${\rm log}\,\sigma_{i}^{2}$ are learned from two GCNs with shared first-layer training parameters. Samples of $q(Z|X,A)$ can be obtained by using the reparameterization trick [58].
The decoder relies on the inner product of the latent representations of each pair of nodes. In the case of GAE, the reconstruction of the adjacency matrix ${A}$ is formulated as follows:
@@
where $\sigma$ is the sigmoid function. The decoder for VGAE is probabilistic and can be expressed as follows:
@@
The training objective of VGAE is to recover information about the graph from the embedding matrix $Z$. Therefore, our objective is to estimate the model parameters by maximizing the logarithm of the marginal distribution, which can be lower-bounded by ELBO:
@@
$=$ $\mathbb{E}_{Z\sim q_{\theta}(Z|X,A)}\log p(A|Z)$
$-D_{KL}(q_{\theta}(Z|X,A)||p(Z)).$ (10)
Here, the inequality is derived from Jensen’s inequality [59], and $D_{KL}(\cdot||\cdot)$ represents the Kullback-Leibler (KL) divergence [60]. We set $p(Z)$ to be a Gaussian normal distribution, i.e.,
@@
III-B Problem Formulation
For (variational) graph auto-encoder, the objective is to learn a graph encoder $f_{\theta}(\cdot)$ with parameters $\theta$ that outputs interpretable low-dimensional stochastic latent variables $Z=f_{\theta}(G)\in\mathbb{R}^{N\times d}$ ($d\ll n$). Simultaneously, such learned node representations are capable of tackling a wide variety of tasks related to nodes.
In our study, we aim to improve the link prediction performance in (variational) graph auto-encoder by improving its overall framework, aligning with the experiments conducted in [12]. To achieve this, we will learn disentangled node representations using the disentangled graph encoder $f_{\theta}$.
For each node $u$, the output $z_{u}$ from the disentangled graph encoder will be a disentangled representation. Specifically, assuming there are $K$ latent factors to be disentangled, $z_{u}$ is composed of $K$ independent components, i.e., $z_{u}=\left[z_{u,1},z_{u,2},\dots,z_{u,K}\right]$, where $z_{i,k}\in\mathbb{R}^{\frac{d}{K}}$ ($1\leq k\leq K$). The $k$-th component $z_{u,k}$ is used to characterize aspects of node $u$ that are related to the $k$-the latent factor. In this context, we assume the underlying factors generating observed graph-structured data are mutually independent which is well accepted in the field of disentangled representation learning [20, 40]. Our goal is to learn factorized and expressive representations where each component is independent and exclusively corresponds to a single ground-truth factor behind the graph formulation. These representations will enhance performance in widely-used graph analytic tasks.
Figure 2: The whole framework of the proposed DVGA. It takes in nodes and their neighbors along with their feature vectors. The dynamic disentangled encoder utilizes DGCN, which employs a dynamic assignment mechanism that considers neighborhoods induced by $K$ distinct factors and neighborhoods in the $k$-th component space. After channel-wise convolution, the resulting component representations are merged and fed into flows for learning expressive disentangled node representations, which are then sent into a factor-wise decoder to reconstruct the adjacency matrix. A classifier $g_{\phi}$ is employed to promote the independence of component representations among several latent factors. The joint optimization of the factor-wise auto-encoder and independence regularization enhances the disentanglement. In this example, we assume the existence of three latent factors, each of which corresponds to one of the three channels.
IV Proposed Mechanism
@@
IV-A Dynamic Disentangled Graph Encoder
Here, we introduce the disentangled graph encoder that can learn disentangled node embeddings. It will consist of a disentangled graph convolutional neural network, which we denote as DGCN. We assume that each node consists of $K$ independent components, indicating the presence of K latent factors that need to be disentangled. For a node $u\in V$, its hidden representation $z_{u}$ is expressed as $\left[z_{u,1},z_{u,2},\dots,z_{u,K}\right]$, where $z_{u,k}\in\mathbb{R}^{\frac{d}{K}}$ ($1\leq k\leq K$) is used to describe $k$-th aspect of node $u$.
@@ IV-A Dynamic Disentangled Graph Encoder
where $W_{k}\in\mathbb{R}^{n\times\frac{d}{K}}$ and $b_{k}\in\mathbb{R}^{\frac{d}{K}}$ are the parameters of $k$-th subspace. For a comprehensive understanding of aspect $k$ for node $u$, we need to extract information from its neighborhood. This involves constructing $z_{u,k}$ using both $c_{u,k}$ and $\{c_{v,k}:(u,v)\in E\}$. So we provide three hypotheses for the relationships between nodes:
Hypothesis 1: The generation of links between nodes results from the interaction of various hidden factors.
Hypothesis 2: If there is similarity in the $k$-th component space between node $u$ and its neighboring node $v$, it indicates that the link between these two nodes may be attributed to factor $k$.
@@
Hypothesis 1 aligns with our realistic understanding of the reasons behind the formation of graphs. It suggests that the links between nodes can be influenced by various factors. In comparison to previous research on disentanglement, this hypothesis is closer to real-world situations [47, 45]. For example, in a friendship network, the acquaintance between two individuals may result from relationships such as colleagues, classmates, or shared interests. Hypotheses 2 and 3 correspond to the first-order and second-order proximities within the topological structure influenced by factor $k$, respectively [25]. This insight inspires us to explore $K$ subspaces to identify latent factors.
Building upon these three hypotheses, we can learn the disentangled representations of nodes by leveraging dynamic assignment mechanism [21], which is a disentangle layer:
@@
where iteration $t=1,2,\dots,T$. $p_{u,v}^{k}$ represents the probability of the connection between nodes $v$ and $u$ attributed to the $k$-th factor, while $q_{v,u}^{k}$ denotes the importance of aggregating node $v$ with respect to node $u$ in the $k$-th component space. They satisfies $p_{u,v}^{k}\geq 0$, $q_{v,u}^{k}\geq 0$, $\sum_{k=1}^{K}p_{u,v}^{k}=1$ and $\sum_{k=1}^{K}q_{v,u}^{k}=1$. $\alpha$ and $\beta$ serve as trade-off coefficients and we treat them as trainable parameters and adjust their range to be within $(0,1)$. The dynamic assignment mechanism will iteratively infer $p_{u,v}^{k}$ ,$q_{v,u}^{k}$ and construct $z_{u,k}$. Meanwhile, this mechanism enables the coexistence of multiple relationships between nodes. It’s important to note that there are a total of $L$ disentangle layers, and in each layer $l\leq L-1$, the value of $c_{u,k}$ is ultimately assigned as $z_{u,k}^{T}$.
@@
By applying Eqs. (2), (3), (4) and replacing GCN with DGCN, we finally obtain the dynamic disentangled graph encoder.
In contrast to existing holistic graph encoders, this dynamic disentangled graph encoder comprises $K$ channels, enabling the identification of complex heterogeneous latent factors and the depiction of multiple relation aspects.
IV-B Enhancing Posterior Flexibility through Component-wise Normalizing Flows
It is evident that when the inference process depends exclusively on the dynamic disentangled graph encoder, DVGA will generally provide a distribution that is factorized. This suggests that the individual component representation $z_{u,k}$ is not only independent of each other but also mutually independent for each $d$-dimensional vector. It is unnecessary and undesirable to have the latter because we aim for each $d$-dimensional component to effectively capture sufficient information about each factor. In this section, we aim to improve the flexibility of each $d$-dimensional component representation so that it can effectively capture the complex properties of the factors.
Normalizing Flows (NF) [61] are commonly employed to expand the range of posterior distribution families. They impose explicit density functions for the mixing distributions in the hierarchy[14, 38]. To better learn the representation of each factor, we introduce a flow-based, component-wise distribution as the posterior distribution for the latent variables in DVGA.
The main flow process involves applying separate flows to the $K$ components of $z_{u}$, with each flow consisting of $M$ steps (e.g., $M$ = 4). Each step includes an affine coupling layer, which guarantees that the latent variable $z_{u,k}$ follows a flexible distribution. More precisely, we follow the following iteration process to apply flows to each component $1\leq k\leq K$, denoted as $m=1,2,\dots,M$:
@@ IV-B Enhancing Posterior Flexibility through Component-wise Normalizing Flows
Here, $s_{k}(\cdot)$ and $t_{k}(\cdot)$ denote transformation functions (mapping the input from $d^{\prime}$ to $d-d^{\prime}$ dimension space), $s(z_{u,k,1:d^{\prime}}^{(m-1)})=\left[s_{1}(z_{u,k,1:d^{\prime}}^{(m-1)}),s_{2}(z_{u,k,1:d^{\prime}}^{(m-1)}),\dots,s_{d-d^{\prime}}(z_{u,k,1:d^{\prime}}^{(m-1)})\right]^{\top}$, $t(z_{u,k,1:d^{\prime}}^{(m-1)})=\left[t_{1}(z_{u,k,1:d^{\prime}}^{(m-1)}),t_{2}(z_{u,k,1:d^{\prime}}^{(m-1)}),\dots,t_{d-d^{\prime}}(z_{u,k,1:d^{\prime}}^{(m-1)})\right]^{\top}$. The $\odot$ represents the Hadamard product. The output of the affine coupling layer is obtained by concatenating $z_{u,k,1:d^{\prime}}^{(m)}$ and $z_{u,k,d^{\prime}+1:d}^{(m)}$. We swap or randomly shuffle the dimensions to enhance the mixing of information between two affine coupling layers.
@@
IV-C Factor-wise Decoder
Unlike the traditional decoder design found in the (variational) graph auto-encoder and other approaches that broadly utilize graph information to design decoder [14, 15, 17], both DGA and DVGA utilize a novel factor-wise decoder to predict edge connections between nodes. This architecture promotes link prediction performance by encouraging the decoder to integrate information at the factor level.
As we already know, the formation of real-world graphs is typically influenced by multiple latent heterogeneous factors. Therefore, if any factor contributes to the presence of a connection between two nodes, an edge will be established between them in the final outcome. Given this insight, we make alterations to the conventional decoder. The structure of the proposed decoder is also illustrated in Figure 2. First, we calculate the cosine similarity between each pair of nodes $z_{u}=\left[z_{u,1},...,z_{u,K}\right]^{\top}$ and $z_{v}=\left[z_{v,1},...,z_{v,K}\right]^{\top}$ for each of the $K$ components. The output vector $A_{factor}$ is a three-dimensional matrix in $\mathbb{R}^{K\times N\times N}$. It represents the predicted connection status of nodes in $K$ component spaces that are related with $K$ factors. Each component space link yields a $N\times N$ matrix as the outcome. This vector $A_{factor}$ is then fused by performing a max-pooling operation to the prediction of each two-node connection of the $K$ channels. This process results in the final state of the edge connection. The entire block, denoted as $\mathcal{D}$, is ultimately merged with inner products constructed from $Z$. The additional inner product link operates similarly to the residual connections utilized in residual networks [62]. Finally, the output of the factor-wise decoder is as follows: